You are a financial advisor. You have a client who wants to save $2,000,000 for retirement...

  1. You are a financial advisor. You have a client who wants to save $2,000,000 for retirement at age 70. The client is currently 25.
  1. What monthly payment does the client need to make if the investment earns 6% interest per annum and the payment is made at the beginning of each month.  Write the formula and the value in the space for the answer to this question.
  2. What is the payment if the client waits until age 40 to start saving, the investment earns 5% per annum and the payment is made at the end of each month?  Write the formula and the value in the space for answer to this question.                 
Answer #1

FV =2000000
Number of months =(70-25)*12 =420

a. Beginning monthly payment =FV/((1+r)*((1+r)^n-1)/r) =2000000/((1+6%/12)*((1+6%/12)^420-1)/(6%/12))=1396.81

b. Number of months =(70-40)*12 =360
Monthly payment at end of month =FV/((1+r)^n-1)/r) =2000000/(((1+5%/12)^360-1)/(5%/12))=2403.10